OpenStudy (anonymous):

If 200 g of iron absorbs 22,000 cal of heat, what will be the change in temperature (specific heat of iron =0.11 cal/g°C).

OpenStudy (perl):

we can use the equation Q = mass * specific heat * change in Temperature

OpenStudy (anonymous):

So, could you help me work this exact problem?

OpenStudy (perl):

We can plug into the formula above 22000 = 200 * (0.11) * delta T and solve for delta T

OpenStudy (anonymous):

Okay, so then we would have to do the opposite and divide correct? so it would be 22000/ 200(0.11) = delta T?

OpenStudy (anonymous):

Sorry, I'm not very good at this stuff

OpenStudy (perl):

yes thats correct

OpenStudy (perl):

just be careful with the parentheses

OpenStudy (perl):

22000/ ( 200(0.11) ) = delta T

OpenStudy (anonymous):

so typing that into my TI83, I would type it exactly as you just listed it?

OpenStudy (perl):

yes you can do that. or 22000 / ( 200 * 0.11)

OpenStudy (anonymous):

I got 1000, is that correct?

OpenStudy (perl):

yes

OpenStudy (perl):

so there will be a change of 1000 degrees Celsius temperature

OpenStudy (anonymous):

AWESOME, thank you your great